Question 4

Corporate income taxes are 18 percent and 35 percent in Singapore and India, respectively. Assume the nations have a tax treaty that eliminates double taxation by allowing a tax credit for foreign taxes.
What is the effective tax rate for a Singaporean firm doing business in India? (use pp.pp%)
Answer #1

effective tax rate will be 35%

Assume a business income of $100. The firm will pay 18% tax in India = $18

In Singapore, the firm will have to pay 35%*100 -18 = $17

Total tax paid = 18+17=35

Hence tax rate = 35%